# MaeFuse: Transferring Omni Features with Pretrained Masked Autoencoders for Infrared and Visible Image Fusion via Guided Training

### v1

@@ 3.1 Network Architecture
The proposed MaeFuse primarily consists of two architectures: 1) A pretrained encoder and decoder based on MAE, utilizing the encoder weights from MAE to directly obtain our feature vectors. 2) A learnable two-layer fusion network. We illustrate the proposed method in Figure 2.
@@
Figure 3: The schematic diagram illustrates our two-stage training process. The left side of the fusion module displays the original features outputted by the encoder, whereas the right side shows the features outputted by the fusion layer. The first stage involves aligning the feature domains of the fusion layer and the encoder. The second stage progresses with training using a fusion loss function Eq. (4). This two-stage training strategy is designed to effectively circumvent the issue of becoming trapped in local optima.
@@
Fusion Layer. In the fusion process, our objective is for the network to selectively integrate modal information in various areas, drawing on the feature information from both modalities. Essentially, the aim is to ensure that the fusion outcome in each region is enriched with a higher density of information. The initial layer of the fusion layer, known as the Comparative Fusion Module (CFM), primarily enables the cross-learning of feature information from modalities $\Phi_{V}$ and $\Phi_{I}$. The output of CFM, represented by $\Phi_{D}$, predominantly captures essential contour information from both modalities. However, it is important to note that during the cross-learning process, this module often loses a significant amount of detail information.
To mitigate the problem of the Comparative Fusion Module (CFM)’s output lacking fine details, we incorporated the Merging Fusion Module (MFM). Within this framework, the feature vector $\Phi_{D}$, produced by the CFM, acts as a guide for contour features, aiding in the re-fusion of the initial encoded features. The original encoded features $\Phi_{V}$ and $\Phi_{I}$ from the two modalities are then compared with $\Phi_{D}$. In regions where the original encoding closely aligns with $\Phi_{D}$, the information is retained, while dissimilar regions are de-emphasized. This method ensures that the resulting fusion feature vector not only retains the contour information of $\Phi_{D}$ but also enriches it with more detailed information.
Mathmatically, CFM is denoted as $\mathbb{C}(\cdot)$, and MFM as $\mathbb{M}(\cdot)$. $\Phi_{D}$ represents the output content of CFM, while $\Phi_{F}$ is the final output result. Thus, we have
@@
3.2 Loss Function
For the purpose of maintaining image information integrity, our approach predominantly incorporates the $\mathcal{L}_{1}$ loss function along with the gradient loss:
$\mathcal{L}_{\text{total}}$ $=\mathcal{L}_{\text{int}}+\alpha\mathcal{L}_{\text{grad}},$
@@ 3.2 Loss Function
$\mathcal{L}_{\text{grad}}$ $=\frac{1}{HW}\left\|\left|\nabla F\right|-\max\left(\left|\nabla V\right|,\left|\nabla I\right|\right)\right\|.$ (4)
@@
Figure 4: Visual comparisons utilizing high-exposure images on the MSRS dataset. Compare these with both state-of-the-art deep learning-based fusion methods and downstream task-driven fusion methods
3.3 Guided Training Strategy
@@
Algorithm 1 Two-Stage Training
@@
3.3.1 Two-Stage Training
The MAE encoder can represent image data in a feature domain, where we need to fuse features of two modalities. Therefore, we first need to align the feature domain output by the fusion layer with the feature domain of the encoder. This alignment is our goal for guided learning. Once our feature domains are aligned, we can use a fusion loss function to promote the integration of information from the two modalities. Ignoring this phased training approach may lead to getting trapped in local optima. Our training strategy schematic is shown in Figure 3.
@@ 3.3.1 Two-Stage Training
Figure 5: Visualization results of the training process. The first and second rows represent the training process of CFM and MFM, respectively. When conducting training for each module, there are two steps involved: modality alignment training based on the average of encoder and training based on the fusion loss. These two steps are known as two-stage training.
3.3.2 Hierarchical Training of Network Structures
@@
For the CFM and MFM, we adopt a sequential guided training approach. Initially, we focus on the CFM, using a two-stage training strategy. This involves directly connecting the CFM’s output to the decoder while keeping the MFM inactive. This step prevents the direct connection of the MFM with the encoder from diminishing the cross-learning effect in CFM. Once CFM training is complete, we lock its weights. Subsequently, we apply a similar two-stage strategy for the MFM, where its input is the output of the now-trained CFM, and its output is linked to the decoder. This approach ensures that the MFM effectively learns the fusion features from the CFM, thereby maximizing the retention of information acquired by the CFM.
Method MSRS LLVIP RoadScene TNO
EN CC SCD EN CC SCD EN CC SCD EN CC SCD
DeFusion 6.342 0.596 1.291 7.141 0.645 1.299 6.857 0.645 1.27 6.683 0.553 1.491
ReCoNet 5.092 0.556 1.255 6.186 0.662 1.472 7.086 0.65 1.636 6.711 0.543 1.681
SuperFusion 6.587 0.601 1.657 7.272 0.662 1.533 7.000 0.621 1.375 6.886 0.527 1.579
MetaFusion 6.365 0.623 1.51 7.101 0.638 1.379 7.407 0.647 1.595 7.237 0.542 1.712
BDLFusion 6.093 0.653 1.563 7.165 0.644 1.371 7.01 0.646 1.339 6.968 0.549 1.632
MaeFuse 6.575 0.628 1.674 7.362 0.672 1.59 7.081 0.672 1.644 6.978 0.576 1.730
Table 1: Quantitative results of the IVF task on four datasets. Red and blue represent the best and second-best values, respectively.
@@
Figure 7: Visual comparison for ‘010418’ in LLVIP IVF dataset.
4 Experiments
4.1 Implementation Details
We conducted experiments on four representative datasets (TNO Toet and Hogervorst (2012), LLVIP Jia et al. (2021), RoadScene Xu et al. (2020b) and MSRS Tang et al. (2022c)). The AdamW optimizer was employed to update parameters of different modules of the model. We utilized PyTorch’s automatic mixed precision (AMP) for computation. And we trained the network structure using a fixed learning rate of $1e^{-4}$ All experiments were implemented under the PyTorch framework.
4.2 Evaluation in Multi-modality Image Fusion
We conducted qualitative and quantitative analyses with eight state-of-the-art competitors, including ReCoNet Huang et al. (2022), SuperFusion Tang et al. (2022a), DeFusion Liang et al. (2022), SwinFusion Ma et al. (2022), CDDFuse Zhao et al. (2023b), SeAFusion Tang et al. (2022b), MetaFusion Zhao et al. (2023a), and BDLFusion Liu et al. (2023c).
Qualitative comparisons. To demonstrate that the MAE pretrained encoder can effectively extract high-level visual information, we categorize the comparative methods into two groups: those without fusion training driven by downstream tasks and those with fusion training driven by downstream tasks. In the exposed images of the MSRS dataset, we find that methods not driven by downstream tasks fail to effectively represent object information and are easily influenced by high exposure. However, methods driven by downstream tasks can to some extent reveal object information. Our method, not driven by downstream tasks, also shows equally good or even better results, indicating that our encoder can extract and fuse high-level visual information. This is exemplified in Figure 4.
Since our encoder can establish a unity between low-level and high-level visual information, the fused images by MaeFuse also maintain texture details very well. In the data of LLVIP, our fused images fully preserve the texture information of the images, such as the grids in the picture. Moreover, our fused images are not influenced by infrared images, thus avoiding artifacts. For example, we can still clearly see the text on buildings. Our fused images only highlight significant infrared information on people, which demonstrates our model’s ability to extract high-level visual information. The content is shown in Figure 7.
Quantitative comparison. We also report in Table 1 the numerical comparison results of our method against five other fusion competitors on four datasets: MSRS, LLVIP, RoadScene, and TNO. We use three objective metrics, EN, CC, and SCD, for analysis, which means that the fused images are compared in terms of the amount of graphic information, correlation with the input images, color distribution, and multi-scale structural similarity. It is clearly observed that our method demonstrates superiority in these statistical metrics.
4.3 Ablation studies
@@
To demonstrate the effectiveness of our guided training strategy, we conducted ablation experiments on both two-stage training and hierarchical training strategies. First, for the two-stage training, we trained a group of fusion layers directly with the fusion loss function for 50 iterations, while another group of fusion layers underwent domain alignment training for 25 iterations, followed by another 25 iterations of training with the same fusion loss function. From the qualitative results, we observed that the data not undergoing two-stage training gradually failed to express image information, falling into local traps. In contrast, data with two-stage training rapidly aligned feature domains in the first 25 iterations, then refined image details with the fusion loss function in the latter 25 iterations, thus progressively eliminating visual issues caused by the ViT block effect. From a quantitative standpoint, the loss results without two-stage training gradually stabilized, indicating a local optimum trap. Meanwhile, data with two-stage training initially did not move towards the best fusion loss result but rapidly converged after domain alignment and achieved better outcomes, also showing a continuous downward trend. These points prove the effectiveness of our training method. Qualitative results are shown in Figure 6, and quantitative results are presented in Figure 8.
Figure 8: The training process for two-stage training. Blue represents two-stage training, with the first 25 iterations as the first stage and the latter 25 iterations as the second stage. Orange represents training without using the two-stage approach. The left graph shows the gradient loss in fusion, and the right graph shows the $\mathcal{L}_{1}$ loss.
Figure 9: Visualization results of hierarchical training ablation experiment. The first row shows the results with activated CFM weights, and the second row shows the results with locked CFM weights. The number of iterations corresponding to each image increases from left to right.
@@
To demonstrate the effectiveness of hierarchical training, we conducted an experiment using the same fusion training function. In this experiment, one group had the weights of the CFM activated, while the other group had the CFM weights locked. As shown in the results, the training with locked weights better maintained the original input features. In contrast, training with activated weights tended to cause the loss of original learning parameters, leading to overfitting of the loss function in the image results. Therefore, hierarchical training is beneficial for learning features content layer by layer. Qualitative results are shown in Figure 9.
4.4 Discovery
Infrared and visible image fusion tasks often operate under the assumption that infrared images lack detail. Contrary to this belief, our experiments reveal that infrared images actually possess significant detail. To uncover these hidden details, we employ Gamma correction to the infrared image, some results are shown in Figure 10.
Figure 10: Lighting the infrared image by Gamma correction. With the decrease of $\gamma$, more and more details are being revealed.
We can clearly see that infrared images have relatively complete image contour information. At this point, we wanted to know if the information from the infrared images has a strong effect. We used the feature vector of the infrared image as the input for the MFM module, replacing the CFM module. Figure 11 shows the results and demonstrate outstanding performance. When the lighting conditions are normal, the performance does not decrease, and it can improve performance in extreme overexposure. Therefore, how to better utilize the information from infrared images can be a consideration of our future work!
Figure 11: The fused results by replacing the CFM with infrared features under normal (the first two rows) and overexposure (the last two rows) light conditions.
5 Conclusion
This paper presents relevant experiments demonstrating the effectiveness of using a pretrained encoder with Masked Autoencoders (MAE). This approach enables the encoded features to capture both high-level and low-level visual information efficiently. Additionally, our guided training strategy significantly accelerates the network’s convergence towards the optimal solution, while minimizing the risk of entrapment in local optima. A notable discovery is the abundance of contour information in infrared images. Identifying strategies for the effective utilization of this information represents a promising avenue for future fusion research.

### v2

@@ III-B Network Architecture
The proposed MaeFuse primarily consists of two architectures: 1) A pretrained encoder and decoder based on MAE, utilizing the encoder weights from MAE to directly obtain our feature vectors. 2) A learnable two-layer fusion network. We illustrate the proposed method in Fig. 3.
MAE. Since MAE is a type of Vision Transformer (ViT), we first briefly review the overview of ViT. For a 2D image $I$, ViT represents $I$ as a sequence of 1D patch embeddings $T=\{t_{i}\mid i=1,\ldots,n\}$, where $n$ denotes the total number of patches. Each patch embedding $t_{i}$ corresponds to a fixed-size section of $I$. Additionally, positional embeddings are added to these patch embeddings to preserve the positional information of the patches. Besides positional embeddings, there is also a classification patch embedding, $[CLS]$, included as an additional learnable embedding, which serves as the global image representation. The complete set of embeddings is referred to as a set of tokens $T=\{t_{i}\mid i=CLS,1,\ldots,n\}$.
The tokens $T$ are then passed through the Transformer encoder, which is composed of multiple Transformer layers stacked together. Each Transformer layer contains Layer Normalization (LN) layers, Multi-Head self-Attention (MHA) modules, and Multi-Layer Perceptron (MLP) blocks. The output of the tokens from the $(l+1)$-th Transformer layer can be expressed as:
$T^{l+1}$ $=\text{MHA}(\text{LN}(T^{l}))+T^{l},$
$T^{l+1}$ $=\text{MLP}(\text{LN}(T^{l+1}))+T^{l+1}.$ (6)
Specifically, within the multi-head self-attention (MHA) block, tokens are transformed into queries, keys, and values:
$Q^{l+1}$ $=T^{l}\cdot W_{q}^{l+1},$
$K^{l+1}$ $=T^{l}\cdot W_{k}^{l+1},$
$V^{l+1}$ $=T^{l}\cdot W_{v}^{l+1},$ (7)
where $W_{q}^{l+1}$, $W_{k}^{l+1}$, and $W_{v}^{l+1}$ are learnable weights.
MAE is based on this structure and conducts self-supervised training. Specifically, we randomly select some tokens to mask $M=\{m_{i}\mid m_{i}\in\{0,1\},i=1,\ldots,n\}$, and then use the unmasked parts to predict the masked areas. Since we need to recover the entire image from only 25% of the remaining regions, we must use a small amount of information to acquire semantic information to guide the recovery of the masked areas. This allows our encoder to obtain good high-level semantic information. Additionally, in the image recovery process, we need to match the colors and textures with the remaining parts of the image, so the encoder also acquires valuable low-level visual information. Therefore, good omni features that facilate the low-level and high-level task simulantoursly can be obtained through self-supervised training with random masking.
@@
Fusion Layer. In the fusion process, our objective is for the network to selectively integrate modal information in various areas, drawing on the feature information from both modalities. Essentially, the aim is to ensure that the fusion outcome in each region is enriched with a higher density of information. The first part of the fusion layer is the Comparative Fusion Module (CFM), which primarily facilitates the interactive learning of features from two modalities through two symmetric cross-attention networks. The output of CFM, represented by $\Phi_{D}$, predominantly captures essential contour information from both modalities. However, it is important to note that during the cross-learning process, this module often loses a significant amount of detail information.
To mitigate the problem of the Comparative Fusion Module (CFM)’s output lacking fine details, we incorporated the Merging Fusion Module (MFM). Within this framework, the feature vector $\Phi_{D}$, produced by the CFM, acts as a guide for contour features, aiding in the re-fusion of the initial encoded features. The original encoded features $\Phi_{V}$ and $\Phi_{I}$ from the two modalities are then compared with $\Phi_{D}$. In regions where the original encoding closely aligns with $\Phi_{D}$, the information is retained, while dissimilar regions are de-emphasized. This method ensures that the resulting fusion feature vector not only retains the contour information of $\Phi_{D}$ but also enriches it with more detailed information. And the FFN structure has strong non-linear learning capabilities, allowing us to better align our fusion domain with the domain of the decoder.
Mathematically, CFM is denoted as $\mathbb{C}(\cdot)$, MFM as $\mathbb{M}(\cdot)$ and FFN as $\mathbb{FFN}(\cdot)$. $\Phi_{D}$ represents the output content of CFM, $\Phi_{M}$ represents the output content of MFM, while $\Phi_{F}$ is the final output result. Thus, we have
@@
Here, we have removed the mask from MAE, and we directly use the $\mathcal{L}_{1}$ loss function to guide the decoder in reconstructing the original input image.
Figure 4: The image ‘00909N’ scene is from the MSRS dataset. The first row shows visible images, while the second row shows infrared images. The first column contains the original images, the second column contains gradient images, and the third column contains second derivative images. $\nabla$ is the Sobel operator, and $\Delta$ is the Laplacian operator.
III-C Loss Function
Firstly, we need to train the decoder to acquire the ability to recover the original image from image features:
$\mathcal{L}_{\text{decoder}}$ $=\left\|\mathbb{D}(\mathbb{E}(I))-I\right\|.$ (11)
Here, we lock the encoder $\mathbb{E}(\cdot)$ and then apply $\mathcal{L}_{1}$ loss between the decoder’s $\mathbb{D}(\cdot)$ output image and the original image to update the decoder’s weight information.
In our previous formula derivation 5, we know that if our fusion results benefit downstream tasks, then our encoder needs to acquire sufficient semantic information, and our fusion layer also needs to reasonably learn and utilize this information. For the encoder, we addressed the issue by utilizing the pretrained MAE encoder. As for the fusion layer, we use texture loss to guide the network to retain semantic information. The gradient information obtained from the first-order derivative of the image also contains semantic texture information to some extent, so our loss function uses the $\mathcal{L}_{1}$ loss function, the gradient loss function along with the laplacian loss:
$\mathcal{L}_{\text{total}}$ $=\mathcal{L}_{\text{int}}+\alpha\mathcal{L}_{\text{grad}}+\beta\mathcal{L}_{\text{laplacian}},$
@@ III-C Loss Function
$\mathcal{L}_{\text{grad}}$ $=\frac{1}{HW}\left\|\left|\nabla F\right|-\max\left(\left|\nabla V\right|,\left|\nabla I\right|\right)\right\|,$
$\mathcal{L}_{\text{laplacian}}$ $=\frac{1}{HW}\left\|\left|\Delta F\right|-\max\left(\left|\Delta V\right|,\left|\Delta I\right|\right)\right\|.$ (12)
Generally, fusion work commonly uses gradients to represent texture information, as proposed by the GTF [64] method, which has become a standard loss function in fusion. However, these texture information might include unwanted elements (such as overexposure, smoke). Here, we explored the effect of second derivatives on fusion. Since object information generally has distinct boundaries, the gradients can change dramatically; whereas for overexposure and similar effects, there is a gradual change, and the magnitude of gradient changes are not as severe. After calculating the second derivatives, these values will be very small, allowing us to reduce the interference from non-object contour information to some extent. As shown in Fig. 4, combining the use of gradients and second derivatives can achieve a better fusion effect.
Figure 5: The schematic diagram illustrates our two-stage training approach. The left of each fusion module displays the features obtained by the encoder, whereas the right shows the features obtained by the fusion layer. The first stage involves aligning the feature domains of the fusion layer and the encoder. The second stage progresses with training using a fusion loss function Eq. (12). This two-stage training strategy is designed to effectively circumvent the issue of becoming trapped in local optima.
@@
By using the alignment function mentioned above, we can ensure that the current results are comparable to the mean fusion, thus aligning the fusion domain with the encoder’s feature space. On this basis, we can then perform fusion training to achieve a good result.
III-D Guided Training Strategy
@@
Essentially, our guided training strategy is a variant of knowledge distillation, where we use existing knowledge to quickly teach an uninitialized model structure to learn on a relatively good prior basis. For our fusion work, this is akin to using the encoder’s output to guide the initialization of the fusion layer’s weights. The reason we use the mean fusion of encoder output features as a guide here draws from the idea of residual connections; our initialization is similar to a direct residual connection. As we gradually train the fusion layer, we can ensure that the minimum performance is no worse than the result of mean fusion. Another point is the desire for the fusion layer’s output domain to align with the decoder’s domain, making the fusion training more effective. Broadly speaking, for infrared and visible image fusion, or other tasks without ground truth (GT), we often use an approximate loss function for training. The convergence ability of this type of loss is not as good as that with GT, and it inherently carries some bias. By using guided learning, we can effectively inject certain priors into the model, thereby reducing the final deviation from the actual GT, making it closer to the real GT. Specific steps are detailed in the following two subsections.
III-D1 Two-Stage Training
The MAE encoder can represent image data in a feature domain, where we need to fuse features of two modalities. Therefore, we first need to align the feature domain output by the fusion layer with the feature domain of the encoder. This alignment is our goal for guided learning. Once our feature domains are aligned, we can use a fusion loss function to promote the integration of information from the two modalities. Ignoring this phased training approach may lead to getting trapped in local optima. Our training strategy schematic is shown in Fig. 5.
Algorithm 1 The Proposed Guided Training Strategy
@@ III-D1 Two-Stage Training
Figure 6: Illustration of the proposed guided training process. The first and second rows represent the training process of CFM and MFM, respectively. When conducting training for each module, there are two steps: modality alignment training based on the average of encoder and training based on the fusion loss. These two steps are known as two-stage training.
III-D2 Hierarchical Training of Network Structures
@@
For the CFM and MFM, we adopt a sequential guided training approach. Initially, we focus on the CFM, using a two-stage training strategy. This involves directly connecting the CFM’s output to the decoder while keeping the MFM inactive. This step prevents the direct connection of the MFM with the encoder from diminishing the cross-learning effect in CFM. Once CFM training is complete, we freeze its weights. Subsequently, we apply a similar two-stage strategy for the MFM, where its input is the output of the now-trained CFM, and its output is linked to the decoder. This approach ensures that the MFM effectively learns the fusion features from the CFM, thereby maximizing the retention of information acquired by the CFM.
IV Experiments
IV-A Implementation Details
We conducted experiments on four representative datasets: TNO [65], M3FD [17], RoadScene [66], and MSRS [40]. The AdamW optimizer was utilized to update the parameters of various model modules. For computations, we employed PyTorch’s Automatic Mixed Precision (AMP). The network was trained using a fixed learning rate of $1e^{-4}$. Our hyperparameters were set to $\alpha=1$ and $\beta=2$, and the number of training epochs corresponded to those specified in Algorithm 1. The model was trained on the training set of the MSRS dataset and evaluated on the test sets of MSRS, M3FD, TNO, and RoadScene. All experiments were implemented using the PyTorch framework.
IV-B Evaluation in Multi-modality Image Fusion
We conducted qualitative and quantitative analyses with nine state-of-the-art competitors, including CUFD [67], DIVFusion [35], GAN-FM [68], STDFusionNet [39], NestFuse [34], SeAFusion [15], TarDAL [17], MetaFusion [16], and PSFusion [19]. Among them, methods CUFD, DIVFusion, GAN-FM, STDFusionNet, and NestFuse do not employ downstream tasks for driving. Methods SeAFusion, TarDAL, MetaFusion and PSFusion are semantically driven by downstream tasks, where SeAFusion and TarDAL utilize image detection, and MetaFusion and PSFusion employ image segmentation. Our method does not use downstream tasks for driving.
Figure 7: Qualitative comparison on the ‘00037N’ scene from the MSRS dataset.
Figure 8: Qualitative comparison on the ‘00909N’ scene from the MSRS dataset.
Qualitative comparisons. To demonstrate that the MAE pretrained encoder can effectively extract high-level visual information, we categorize the comparative methods into two groups: those without fusion training driven by downstream tasks (c,d,e,f,g) and those with fusion training driven by downstream tasks (h,i,j,k). In the exposed images of the MSRS dataset, we find that methods not driven by downstream tasks fail to effectively represent object information and are easily influenced by high exposure. However, methods driven by downstream tasks can to some extent reveal object information. Our method, not driven by downstream tasks, also shows equally good or even better results, indicating that our encoder can extract and fuse high-level visual information. This is exemplified in Fig. 7 and Fig. 8. Similarly, when testing on the TNO dataset, we find that MaeFuse performs better than other works even in conditions with fog obstruction. It is able to maximize the visibility of soldiers behind the fog while retaining the surrounding information, as shown in Fig. 9.
Figure 9: Qualitative comparison on the ‘soldier_behind_smoke_1’ scene from the TNO dataset.
Figure 10: Qualitative comparison on the ‘00922’ scene from the M3FD dataset.
Figure 11: Qualitative comparison on the ‘FLIR_04269’ scene from the RoadScene dataset.
Since our encoder establishes a unity between low-level and high-level visual information, the images fused by MaeFuse also maintain texture details well. In the data from M3FD and RoadScene, our fused images fully preserve the texture and color information of the images. Furthermore, our fusion images are not affected by infrared images, thereby avoiding artifacts. Our pictures do not display obvious fusion artifacts, and the final results tend towards a natural image effect. The content is shown in Fig. 10 and Fig. 11.
Overall, for works (c,d,e,f,g) that do not use downstream tasks, there is not a good unification of high-level semantic information and texture information. Works (h,i,j,k) driven by downstream tasks can highlight semantic information under extreme conditions to a certain extent. Among these, we observed that works (h,j,k) driven by segmentation tasks outperform those (i) driven by image detection tasks. Our work, MaeFuse, can easily solve the above issues and achieves results that are on par with or even surpass those driven by segmentation tasks. It is worth mentioning that the PSFusion work uses a custom mask to assist image fusion, which separates object information in the infrared images with the mask, thereby injecting artificial priors into the results. Our work, however, does not use any custom masks for driving, relying solely on the model’s inherent pre-trained weight priors for fusion, which demonstrates our model’s strong learning capability and potential for future improvement!
TABLE I: Quantitative analysis results on RoadScene [66].
Method CC $\uparrow$ SCD $\uparrow$ PSNR $\uparrow$ NAB/F $\downarrow$ NLPD $\downarrow$
CUFD [67] 0.582 1.420 62.23 0.051 0.6797
DIVFusion [35] 0.606 1.524 61.78 0.055 0.6793
GAN-FM [68] 0.621 1.703 62.03 0.046 0.6581
NestFuse [34] 0.635 1.701 61.99 0.031 0.6372
STDFusionNet [39] 0.590 1.447 60.01 0.085 0.7379
SeAFusion [15] 0.614 1.563 61.74 0.063 0.6604
TarDAL [17] 0.528 1.261 59.57 0.352 0.8451
MetaFusion [16] 0.611 1.547 61.81 0.152 0.7348
PSFusion [19] 0.628 1.696 63.10 0.065 0.6842
MaeFuse 0.653 1.714 63.92 0.028 0.5990
TABLE II: Quantitative analysis results on TNO [65].
Method CC $\uparrow$ SCD $\uparrow$ PSNR $\uparrow$ NAB/F $\downarrow$ NLPD $\downarrow$
CUFD [67] 0.455 1.528 61.14 0.072 0.6160
DIVFusion [35] 0.445 1.494 59.98 0.148 0.6799
GAN-FM [68] 0.469 1.666 61.11 0.073 0.5941
NestFuse [34] 0.488 1.726 62.02 0.032 0.5381
STDFusionNet [39] 0.428 1.440 61.46 0.076 0.5958
SeAFusion [15] 0.482 1.728 61.39 0.081 0.5848
TarDAL [17] 0.403 1.343 59.72 0.392 0.7949
MetaFusion [16] 0.488 1.768 61.78 0.208 0.6669
PSFusion [19] 0.498 1.778 61.72 0.092 0.6316
MaeFuse 0.524 1.820 63.48 0.030 0.5272
TABLE III: Quantitative analysis results on M3FD [17].
Method CC $\uparrow$ SCD $\uparrow$ PSNR $\uparrow$ NAB/F $\downarrow$ NLPD $\downarrow$
CUFD [67] 0.463 1.221 61.70 0.027 0.5327
DIVFusion [35] 0.527 1.530 60.41 0.083 0.6152
GAN-FM [68] 0.548 1.700 62.29 0.028 0.5138
NestFuse [34] 0.542 1.578 62.76 0.009 0.4822
STDFusionNet [39] 0.459 1.179 62.40 0.028 0.5047
SeAFusion [15] 0.525 1.586 61.12 0.026 0.5098
TarDAL [17] 0.481 1.443 58.98 0.185 0.6660
MetaFusion [16] 0.555 1.682 62.42 0.147 0.6271
PSFusion [19] 0.559 1.832 60.80 0.086 0.5927
MaeFuse 0.571 1.750 63.66 0.021 0.5028
Quantitative comparison. We also report in Tables III, and III the quantitative comparison results of our method against nine other fusion methods across three datasets (RoadScene, TNO, M3FD). We used five objective metrics, namely CC, SCD, PSNR, Nabf, and NLPD, which are common evaluation indices. Among them, CC measures the linear correlation between two images, used to assess the similarity between images. A high CC value indicates a high degree of similarity. SCD is used to measure the similarity in the frequency domain between two images, applicable for evaluating the preservation of color or spectral information. PSNR is an indicator of image reconstruction quality, commonly used to assess the effects of image compression or restoration. A high PSNR usually indicates lower errors. The NAB/F metric aims to quantify the number of artifacts in an image, used to assess the errors and anomalies introduced during image processing. A lower value indicates better performance. NLPD is used to evaluate the difference in visual quality between two images, especially in the details of multi-scale structures. And a lower value indicates better performance. It is clearly evident that our method demonstrates superiority in these statistical metrics.
IV-C Ablation Study
Figure 12: The training process for two-stage training. The blue line represents two-stage training, with the first 25 iterations as the first stage and the latter 25 iterations as the second stage. The orange line represents training without using the two-stage approach. The left graph shows the gradient loss in fusion, while the right graph shows the $\mathcal{L}_{1}$ loss.
Figure 13: Visualization results of hierarchical training ablation experiment according to different epoches. The first and second row shows the results with activated CFM weights and locked CFM weights, respectively.
@@ IV-C Ablation Study
Figure 15: Visualization results of the average fusion of feature vectors from different layers of the two modalities.
@@
To demonstrate the effectiveness of our guided training strategy, we conducted ablation experiments on both two-stage training and hierarchical training strategies. First, for the two-stage training, we trained a group of fusion layers directly with the fusion loss function for 50 iterations, while another group of fusion layers underwent domain alignment training for 25 iterations, followed by another 25 iterations of training with the same fusion loss function. In these two experiments, the hyperparameters used for the fusion loss functions are consistent with those in the original experiments. From the qualitative results, we observed that the data not undergoing two-stage training gradually failed to express image information, falling into local traps. In contrast, data with two-stage training rapidly aligned feature domains in the first 25 iterations, then refined image details with the fusion loss function in the latter 25 iterations, thus progressively eliminating visual issues caused by the ViT block effect. From a quantitative standpoint, the loss results without two-stage training gradually stabilized, indicating a local optimum trap. Meanwhile, data with two-stage training initially did not move towards the best fusion loss result but rapidly converged after domain alignment and achieved better outcomes, also showing a continuous downward trend. These points prove the effectiveness of our training method. The quantitative results are shown in Fig. 12, while the qualitative results are presented in Fig. 14.
@@
To demonstrate the effectiveness of hierarchical training, we conducted an experiment using the same fusion training function. In this experiment, one group had the weights of the CFM activated, while the other group had the CFM weights locked. As shown in the results, the training with locked weights better maintained the original input features. In contrast, training with activated weights tended to cause the loss of original learning parameters, leading to overfitting of the loss function in the image results. Therefore, hierarchical training is beneficial for learning features content layer by layer. Qualitative results are shown in Fig. 13.
Figure 16: The first and second rows show the ‘00037N’ and ‘00909N’ scenes from the MSRS dataset. The first column contains visible images, the second column contains infrared images, the third column shows the maximum value fusion in the image domain, and the fourth column shows the maximum value fusion in the feature domain.
Figure 17: The results of the qualitative analysis of the segmentation tasks, from top to bottom, every two rows form a comparison group, respectively. The prompts used for Grounded SAM of the first two rows, the second two rows, and the last two rows are ‘Car.’, ‘Person. Building.’, and ‘Soldier.’, respectively.
Since existing methods often utilize downstream tasks to assist with fusion training, our network is able to learn high-level semantic information. However, we have achieved the same or even better results without using downstream tasks. Here, I demonstrate through several ablation experiments that the feature vectors output by our MAE’s pretrained encoder indeed possess high-level semantic information. In Fig. 16, we directly perform maximum value fusion of images from two modalities in the pixel domain, as well as in the feature domain. We find that performing maximum value fusion in the feature space does not result in overexposure, indicating that our feature vectors contain high-level semantic information and have greater weight. In Fig. 15, we visualize the results of averaging the features output from different layers of the MAE encoder. We find that as the number of layers increases, the high-level semantic information in the encoder becomes more apparent, showing that the MAE encoder can indeed capture sufficient high-level semantic information.
IV-D Generalization Ability
To verify that our fused images indeed retain sufficient high-level semantic information and are beneficial for downstream tasks, we conducted the following experiments for testing: semantic segmentation with prompts using Grounded SAM [69], which first uses prompt words to identify the detection target, then uses SAM [70] for semantic segmentation of the corresponding target. We used the most common segmentation methods to perform the segmentation directly and then compared the results. This allows us to fairly demonstrate whether the semantic information of the fused images can be easily recognized by existing segmentation models. If good results are obtained, it indicates that the information is sufficient and can be easily perceived by downstream tasks.
Existing segmentation techniques can perform well under various complex conditions, as shown in Fig. 17:a and Fig. 17:b, where the segmentation results are better than the corresponding visual results in Fig. 8 and Fig. 10. Therefore, if it performs well visually, the downstream tasks will also perform well, which is the core viewpoint of the paper PSFusion [19]. Regarding the details in Fig. 17:a, we can observe that our results do not have segmentation edge errors at the bottom or rear of the car. Similarly, in Fig. 17:b, we can also effectively segment the three distant apartment buildings.
The comparison image in Fig. 17:c is from the TNO [65] dataset, which does not have segmentation labels. Therefore, all works that use downstream tasks for auxiliary training have not seen these images. In the visual results, we can see many images that fail to highlight the soldier from the fog. As a result, many outcomes either do not generate a mask or result in erroneous full-screen segmentation. It is worth mentioning that for TarDAL [17], PSFusion [19] can visually detect the soldier, but the segmentation results show multiple segmentations or missed detections, which we believe indicate poor generalization capabilities. The high-level semantic information in the fused results has a domain gap with that of general natural images, making it difficult to segment the desired results using the powerful SAM. However, our mask successfully obtains the corresponding segmentation results, indicating that our results are beneficial for downstream tasks and possess good generalization capabilities.
V Conclusion
In this paper, we explore the use of feature vectors obtained from the pretrained MAE encoder for image fusion. Using this method, we can provide sufficient high-level semantic information for the fusion results, and it is not necessary to use downstream tasks for auxiliary training. This simplifies the fusion network structure and training method. Additionally, we propose a two-stage training strategy that allows our fusion network’s feature space to quickly align with the feature space represented by the original MAE encoder, solving the issues of limited training data and falling into local optima. Experiments show that using MaeFuse can match or even surpass fusion methods driven by semantic segmentation tasks. Employing Omni features for fusion can become the next direction in fusion work.
Using pretrained networks for fusion, we need to focus on whether the network can simultaneously capture high-level and low-level visual information. It is worth noting that recently, some [71] have also used CycleGAN [72] for fusion with good results, because CycleGAN, when performing domain transformations, needs to acquire both high-level and low-level visual information. From this perspective, the idea of using Omni features for fusion has significant room for development.